# ReferencesSuppReferences

### v3

@@ 2.5 Robustness to Information Leakage
It is interesting to analyze when relative exogeneity may or may not fail due to information leakage about the policy news or some market anticipation of the policy change. Leakage is highly relevant in applications as documented recently in the finance literature. Leakage has the following empirical features. First, it leads to market anticipation of the news that is to be revealed at the event time. This may reduce the variance of $e_{t}$ substantially for $t\in\mathbf{P}$, possibly making the policy shock on the same order of magnitude as the other random variables in the system, failing to dominate them in the event window. Second, leakage can be associated with learning and sluggish market adjustments since the news may initially reach a small number of market participants and, through their reactions, may slowly spread into the market. This is likely to generate serial dependence in the policy shock over the hours or days prior to the scheduled announcement. We introduce these features into the model and discuss how this can alter the identification results. The introduction of leakage into the model depends on the specific context of the event under consideration. Here we focus on FOMC announcements.
@@
where $v_{t}$ is a white noise process with zero mean and unit variance, $|\vartheta_{j}|<\infty$ for all $j=1,\ldots,\,q$ and $q>0$ is a finite integer. Assume that $\sigma_{t}^{2}\rightarrow\infty$ if $t\in\mathbf{P}$ and $|\sigma_{t}^{2}|<\infty$ if $t\notin\mathbf{C}$. The parameter $\phi_{t}$ is the leakage parameter:
@@
Finally, information leakage could also have negative consequences for obtaining good measures of market expectations about the policy news thereby making it difficult to accurately construct the surprise component of the policy action $D_{t}$. Failure to isolate the expected component of the news leads to attenuation bias as asset prices have already responded to the expected part of the policy news. Whether this attenuation bias is important or not depends on how large the information leakage is relative to the unexpected part of the policy news that is driven by $e_{t}$ in the event window. When the leakage is small, one does not need to extend the event window to bracket the pre-release embargo when computing $D_{t}$ since the policy shock during the 30-minute window is still an order of magnitude larger than the other shocks.
@@ 3 Properties of the OLS Event Study Estimator
Let $\sigma_{D}^{2}=\lim_{T_{P}\rightarrow\infty}T_{P}^{-1}\sum_{t=1}^{T_{P}}\mathrm{Var}\left(D_{t}\right)$, $D_{t}^{*}=\sigma_{D}^{-1}D_{t}$ and $Y_{t}^{*}=\sigma_{D}^{-1}Y_{t}$ for all $t\in\mathbf{P}.$ We make the following assumption in order to study the asymptotic properties of $\widehat{\beta}_{\mathrm{ES}}$.
@@
The following theorem establishes the consistency of the OLS event study estimator under relative exogeneity (cf. Assumption 4). Moreover, since $\Delta$ is decreasing in $\sigma_{e,t}^{2}$, the large-sample bias of $\widehat{\beta}_{\mathrm{ES}}$ for estimating the weighted average of standardized MCEs is small when $\sigma_{e,t}^{2}$ is large. For technical reasons inherent to the proof, to establish this result we take the limits as $T_{P}\rightarrow\infty$ and $\min_{t\in\mathbf{P}}\sigma_{e,t}^{2}\rightarrow\infty$ sequentially. This means that $T_{P}\rightarrow\infty$ before $\min_{t\in\mathbf{P}}\sigma_{e,t}^{2}\rightarrow\infty$.
@@ 4.1 Response of Interest Rates to Monetary Policy News
We consider the setting of Nakamura and Steinsson (2018a) which is a special case of (2.1),
$Y_{t}$ $=\beta D_{t}+u_{t},\qquad\qquad t\in\mathbf{P},$ (4.1)
@@
where $Y_{t}$ is the 1-day change in the 2-Year or 5-Year U.S. Treasury instantaneous real forward rate, $D_{t}$ is the policy news shock constructed by the authors as a change over a 30-minute window (see Example 1 above) and $u_{t}$ is an error term. The policy variable can be endogenous, i.e., $\mathbb{E}\left(u_{t}|\,D_{t}\right)\neq 0$. However, Theorem 4 implies that if relative exogeneity (2.4) holds then the OLS event study estimator $\widehat{\beta}_{\mathrm{ES}}$ in (4.1) is consistent for $\beta.$ Moreover, even if $\sigma_{e,t}^{2}$ is finite but much larger than the variance of $u_{t}$, Theorem 3 implies that $\widehat{\beta}_{\mathrm{ES}}$ has low bias.
@@
We propose a simple empirical procedure to check if $\sigma_{e,t}^{2}$ is relatively large enough that $\widehat{\beta}_{\mathrm{ES}}$ has low bias. We apply our analysis to the setting of Nakamura and Steinsson (2018a) for concreteness, noting that it can be straightforwardly applied to any other event study regression. While it is possible to estimate $\mathrm{Var}\left(D_{t}\right)$ using the sample variance of $D_{t}$ in the policy sample, this could be very large even when relative exogeneity is satisfied only for a few announcements since the variance cannot be negative. Additionally, it is more difficult to estimate $\sigma_{u,t}^{2}$ since $u_{t}$ is not observed and the OLS residuals may not be close in probability to the corresponding true errors $u_{t}$ given that $\mathbb{E}\left(u_{t}|\,D_{t}\right)\neq 0$. The key idea to our procedure is that since $u_{t}$ includes macroeconomic news or factors that are present even when there is no announcement, the order of magnitude of its variance can be retrieved from the order of magnitude of the variance of $Y_{t}$ in the control sample $\mathbf{C}$. In fact, in the control sample the variance of $D_{t}$ and $u_{t}$ have the same order of magnitude and given that $\left|\beta\right|<\infty$, the order of magnitude of the variance of $Y_{t}$ for $t\in\mathbf{C}$ is the same as that of $u_{t}$ so long as (4.1) also holds for $t\in\mathbf{C}$. Thus, we can proxy the average variance of $u_{t}$ for $t\in\mathbf{P}$ by the average variance of $Y_{t}$ for $t\in\mathbf{C}$.
@@
We initially simulate the regression (4.1) calibrated to the corresponding regression in Nakamura and Steinsson (2018a) for the control sample, making the draws of $D_{t}$ and $u_{t}$ independent, and in each draw we estimate $\beta$ by OLS. We repeat this many times and compute bias, mean-absolute error (MAE), and mean-squared error (MSE) for this idealized OLS estimator. We name it the “oracle” estimator since it is obtained in the absence of endogeneity and it is efficient by the Gauss-Markov Theorem.
@@
The performance of the oracle estimator is compared to that of the corresponding OLS event study estimator when we allow for correlation between $D_{t}$ and $u_{t}$. In the latter regression, we successively increase the variance of $D_{t}$ and record how much the performance of the event study estimator approaches that of the oracle estimator in the idealized regression. In particular, we determine the threshold value for the variance of $D_{t}$ that ensures that the event study estimator performs as well as the oracle. Lastly, we verify whether the sample variance of $D_{t}$ in Nakamura and Steinsson (2018a) is larger than this threshold. If it is, relative exogeneity likely holds and we can interpret $\sigma_{e,t}^{2}$ as being large enough relative to the variance of $u_{t}$ for the OLS event study estimator to have low enough bias to perform well in practice.
@@
Let $T_{C}$ denote the sample size of the control sample. The treatment sample consists of all regularly scheduled FOMC meeting days from 1/1/2000 to 3/19/2014. The control sample includes all Tuesdays and Wednesdays that are not FOMC meeting days over the same period. This yields $T_{P}=74$ and $T_{C}=762$ when the dependent variable is the 2-Year real yield, and $T_{P}=106$ and $T_{C}=1130$ when it is the 5-Year real yield.
@@
    Estimate the mean, variance and autoregressive coefficient of order 1 from all $Y_{t}$ in the control sample ($t\in\mathbf{C}$). Denote these by $\overline{Y}_{C}$, $\widehat{\sigma}_{Y,C}^{2}$ and $\widehat{\rho}_{Y,C,1}$. Test whether $\widehat{\rho}_{Y,C,1}$ is statistically significant different from zero.
@@
    Repeat 1. for all $D_{t}$ in the control sample ($t\in\mathbf{C}$) to obtain $\overline{D}_{C},$ $\widehat{\sigma}_{D,C}^{2}$ and $\widehat{\rho}_{D,C,1}$.
@@
  4. 4.
    Let $\widehat{\sigma}_{\widetilde{u}}^{2}=\widehat{\sigma}_{Y,P}^{2}\widehat{\sigma}_{D,C}^{2}/\widehat{\sigma}_{D,P}^{2}$ where $\widehat{\sigma}_{Y,P}^{2}$ and $\widehat{\sigma}_{D,P}^{2}$ are the sample variances of $Y_{t}$ and $D_{t}$ in the policy sample, respectively. Note that $\widehat{\sigma}_{\widetilde{u}}^{2}$ is a proxy for the variance of $Y_{t}$ in the control sample. Since $Y_{t}$ is the 1-day change in the forward rate while $D_{t}$ is the change in a 30-minute window, the variance of $D_{t}$ is smaller than that of $Y_{t}$ in the control sample. So we choose $\widehat{\sigma}_{\widetilde{u}}^{2}$ such that the ratio of the variances of $Y_{t}$ in the treatment and control sample is the same as that of $D_{t}$, i.e., $\widehat{\sigma}_{Y,P}^{2}/\widehat{\sigma}_{\widetilde{u}}^{2}=\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}.$ [^19] Note that $\widehat{\sigma}_{\widetilde{u}}^{2}$ approximates the variance of $u_{t}$ for $t\in\mathbf{P}$ when $D_{t}$ and $u_{t}$ are independent and the variance of $u_{t}$ has the same order of magnitude across policy and control sample. If $Y_{t}$ and $D_{t}$ had been constructed as changes over similarly-sized windows so that their variances had similar orders of magnitude, one could instead use $\widehat{\sigma}_{\widetilde{u}}^{2}=\widehat{\sigma}_{Y,C}^{2}-\beta^{2}\widehat{\sigma}_{D,C}^{2}$ with $\beta$ set equal to 0.99, which is the OLS estimate in Nakamura and Steinsson (2018a). [^20] We could choose any other value for $\beta$, the intuition behind the procedure and the results would not change. For $t=1,\ldots,\,T_{P}$ generate $\widetilde{u}_{t}\sim i.i.d.\,\mathscr{N}(0,\,\widehat{\sigma}_{\widetilde{u}}^{2})$ if $\widehat{\rho}_{Y,C,1}$ is not statistically different from zero. Otherwise generate $\widetilde{u}_{t}$ according to $\widetilde{u}_{t}=\widehat{\rho}_{Y,C,1}\widetilde{u}_{t-1}+\widetilde{v}_{t}$ where $\widetilde{v}_{t}\sim i.i.d.\,\mathscr{N}(0,\,(1-\widehat{\rho}_{Y,C,1}^{2})\widehat{\sigma}_{\widetilde{u}}^{2})$ and $\widetilde{u}_{0}=0$.
@@
    Generate $\widetilde{Y}_{t}=\beta\widetilde{D}_{t}+\widetilde{u}_{t}$ for $t=1,\ldots,\,T_{P}$ where $\widetilde{D}_{t}$ is generated in 3., $\widetilde{u}_{t}$ in 4. and $\beta$ is set equal to 0.99 as explained in 4. Run OLS for the regression of $\widetilde{Y}_{t}$ on $\widetilde{D}_{t}$ and label this OLS estimator $\widehat{\beta}_{\mathrm{orcale}}$. Repeat this 5,000 times and compute its bias, MAE and MSE.
@@
    Find the value of $\delta$ such that $\mathrm{MSE}(\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta\right))=\mathrm{MSE}(\widehat{\beta}_{\mathrm{orcale}})$ (or $\mathrm{MAE}(\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta\right))=\mathrm{MAE}(\widehat{\beta}_{\mathrm{orcale}})$) and label it $\delta^{*}\left(\rho\right)$. Define $\widehat{\sigma}_{D,*}^{2}\left(\rho\right)=\widehat{\sigma}_{D,C}^{2}\left(1+\delta^{*}\left(\rho\right)\right).$
  9. 9.
    Repeat 6.-9. for a grid of $\rho$ values in $[-1,1]$.
@@
Steps 1.-9. provide a detailed sensitivity analysis that computes the values of the variance of $D_{t}$ that allows the event study estimator to perform as well as the oracle estimator over a range of values of endogeneity $\rho$, namely $\widehat{\sigma}_{D,*}^{2}\left(\rho\right)$. If $\widehat{\sigma}_{D,P}^{2}\geq\widehat{\sigma}_{D,*}^{2}\left(\rho\right)$ one can be confident that the variance of the policy shock in Nakamura and Steinsson (2018a) is relatively large enough for a degree of endogeneity that satisfies $|\mathrm{Corr}\left(u_{t},\,D_{t}\right)|\leq|\rho|$ for the event study estimator to have low enough bias for its MSE (MAE) to be as low as in the oracle regression with no endogeneity. If one is not interested in the full sensitivity analysis, but rather just in evaluating the values of $\rho$ for which the policy shock in Nakamura and Steinsson (2018a) is large enough to produce a credible event study estimator, one could simply replace $\delta$ in 7. by $\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}$ and instead of 8., find the value of $\rho$ such that $\mathrm{MSE}(\widehat{\beta}_{\mathrm{ES}}(\rho,\,\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}))=\mathrm{MSE}(\widehat{\beta}_{\mathrm{orcale}})$ (or $\mathrm{MAE}(\widehat{\beta}_{\mathrm{ES}}(\rho,\,\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}))=\mathrm{MAE}(\widehat{\beta}_{\mathrm{orcale}})$).
@@
The validity of the sensitivity analysis relies on the following two properties: (i) the model in (4.1) is correct also in the control sample, i.e., for $t\in\mathbf{C}$ and (ii) the order of magnitude of $\widehat{\sigma}_{\widetilde{u}}^{2}$ is a good proxy for the order of magnitude of $T_{P}^{-1}\sum_{t\in\mathbf{P}}\sigma_{u,t}^{2}$. Property (i) is implicitly assumed in typical VAR estimation of simultaneous equations models since typically the relationship between monetary policy and real economic variables is estimated with VARs at monthly, quarterly, yearly frequencies and each observation aggregates both FOMC announcement and non-announcement days. Property (ii) is reasonable since $u_{t}$ for $t\in\mathbf{P}$ represents news and latent factors that are not specific to the announcement and would be present even if there were no FOMC announcement. Thus, its variation is expected to be similar to that in non-announcement days which in turn has the same order of magnitude as that of $Y_{t}$ in non-announcement days.
@@
To provide numerical support for the theoretical results of Theorem 1-4, we look at the change in the bias, MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta\right)$ as $\delta$ increases for a given $\rho\neq 0$ when looking at the 2-Year real yields. The results show that these summary statistics decrease quickly as $\delta$ grows. For example, for a small degree of endogeneity $\rho=0.25$, it is sufficient to increase the variance of the policy variable in the announcement windows nine-fold $\left(\delta=8\right)$ for the event study estimator to exhibit an MSE that is as small as that of the oracle estimator. A similar feature applies to MAE, though a slightly larger $\delta$ is needed. As we raise $\delta$ further, the MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}$ become smaller and smaller relative to those of $\widehat{\beta}_{\mathrm{oracle}}$. In particular, both the MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}$ converge to zero quickly as $\delta$ increase for $\rho=0.25$.
@@
$\widehat{\beta}_{\mathrm{orcale}}$ $\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta=32\right)$ $\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta=40\right)$ $\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta=48\right)$

### v4

@@ 2.5 Robustness to Information Leakage
It is interesting to analyze when relative exogeneity may or may not fail due to information leakage about the policy news or some market anticipation of the policy change. Leakage is highly relevant in applications as documented recently in the finance literature discussed above. Leakage has the following empirical features. First, it leads to market anticipation of the news that is to be revealed at the event time. This may reduce the variance of $e_{t}$ substantially for $t\in\mathbf{P}$, possibly making the policy shock on the same order of magnitude as the other random variables in the system, failing to dominate them in the event window. Second, leakage can be associated with learning and sluggish market adjustments since the news may initially reach a small number of market participants and, through their reactions, may slowly spread into the market. This is likely to generate serial dependence in the policy shock over the hours or days prior to the scheduled announcement. We introduce these features into the model and discuss how this can alter the identification results. The introduction of leakage into the model depends on the specific context of the event under consideration. Here we focus on FOMC announcements.
@@
where $v_{t}$ is a white noise process with zero mean and unit variance, $|\vartheta_{j}|<\infty$ for all $j=1,\ldots,\,q$ and $q>0$ is a finite integer. Assume that $\sigma_{t}^{2}\rightarrow\infty$ if $t\in\mathbf{P}$ and $|\sigma_{t}^{2}|<\infty$ if $t\in\mathbf{C}$. The parameter $\phi_{t}$ is the leakage parameter:
@@
Finally, information leakage could also have negative consequences for obtaining good measures of market expectations about the policy news thereby making it difficult to accurately construct the surprise component of the policy action $D_{t}$. Failure to isolate the expected component of the news leads to attenuation bias as asset prices have already responded to the expected part of the policy news. Whether this attenuation bias is important or not depends on how large the information leakage is relative to the unexpected part of the policy news that is driven by $e_{t}$ in the event window. When the leakage is small, one does not need to extend the event window to bracket the pre-release embargo when computing $D_{t}$ since the policy shock during the 30-minute window is still an order of magnitude larger than the other variables.
@@ 3 Properties of the OLS Event Study Estimator
Let $\sigma_{D}^{2}=\lim_{T_{P}\rightarrow\infty}T_{P}^{-1}\sum_{t=1}^{T_{P}}\mathrm{Var}\left(D_{t}\right)$, $D_{t}^{*}=\sigma_{D}^{-1}D_{t}$, $Y_{t}^{*}=\sigma_{D}^{-1}Y_{t}$ and $Y_{t}^{*}\left(d\right)=\sigma_{D}^{-1}Y_{t}\left(d\right)$ for all $t\in\mathbf{P}.$ We make the following assumption in order to study the asymptotic properties of $\widehat{\beta}_{\mathrm{ES}}$.
@@
The following theorem establishes the consistency of the OLS event study estimator under relative exogeneity (cf. Assumption 4). Moreover, since $|\Delta|$ is decreasing in $\sigma_{e,t}^{2}$, the large-sample bias of $\widehat{\beta}_{\mathrm{ES}}$ for estimating the weighted average of standardized MCEs is small when $\sigma_{e,t}^{2}$ is large. For technical reasons inherent to the proof, to establish this result we take the limits as $T_{P}\rightarrow\infty$ and $\min_{t\in\mathbf{P}}\sigma_{e,t}^{2}\rightarrow\infty$ sequentially.
@@ 4.1 Response of Interest Rates to Monetary Policy News
We consider the regression of Nakamura and Steinsson (2018a) which is a special case of (2.1),
$Y_{t}$ $=\beta D_{t}+\widetilde{u}_{t},\qquad\qquad t\in\mathbf{P},$ (4.1)
@@
where $Y_{t}$ is the 1-day change in the 2-Year or 5-Year U.S. Treasury instantaneous real forward rate, $D_{t}$ is the policy news surprise constructed by the authors as a change over a 30-minute window (see Example 1 above) and $\widetilde{u}_{t}$ is an error term. The policy variable can be endogenous, i.e., $\mathbb{E}\left(\widetilde{u}_{t}|\,D_{t}\right)\neq 0$. However, Theorem 4 implies that if relative exogeneity (2.4) holds then the OLS event study estimator $\widehat{\beta}_{\mathrm{ES}}$ in (4.1) is consistent for $\beta.$ Moreover, even if $\sigma_{e,t}^{2}$ is finite but much larger than the variance of $\widetilde{u}_{t}$, Theorem 3 implies that $\widehat{\beta}_{\mathrm{ES}}$ has low bias.
@@
We propose a simple empirical procedure to check if $\sigma_{e,t}^{2}$ is relatively large enough that $\widehat{\beta}_{\mathrm{ES}}$ has low bias. We apply our analysis to the setting of Nakamura and Steinsson (2018a) for concreteness, noting that it can be straightforwardly applied to any other event study regression. While it is possible to estimate $\mathrm{Var}\left(D_{t}\right)$ using the sample variance of $D_{t}$ in the policy sample, this could be very large even when relative exogeneity is satisfied only for a few announcements since the variance cannot be negative, as discussed above. Additionally, it is more difficult to estimate $\sigma_{\widetilde{u},t}^{2}$ since $\widetilde{u}_{t}$ is not observed and the OLS residuals may not be close in probability to the corresponding true errors $\widetilde{u}_{t}$ given that $\mathbb{E}\left(\widetilde{u}_{t}|\,D_{t}\right)\neq 0$. The key idea to our procedure is that since $\widetilde{u}_{t}$ includes macroeconomic news or factors that are present even when there is no announcement, the order of magnitude of its variance can be retrieved from the order of magnitude of the variance of $Y_{t}$ in the control sample $\mathbf{C}$. In fact, in the control sample the variance of $D_{t}$ and $\widetilde{u}_{t}$ have the same order of magnitude and given that $\left|\beta\right|<\infty$, the order of magnitude of the variance of $Y_{t}$ for $t\in\mathbf{C}$ is the same as that of $\widetilde{u}_{t}$ so long as (4.1) also holds for $t\in\mathbf{C}$. Thus, we can proxy the average variance of $\widetilde{u}_{t}$ for $t\in\mathbf{P}$ by the average variance of $Y_{t}$ for $t\in\mathbf{C}$.
@@
We initially simulate the regression (4.1) calibrated to the corresponding regression in Nakamura and Steinsson (2018a) for the control sample, making the draws of $D_{t}$ and $\widetilde{u}_{t}$ independent, and in each draw we estimate $\beta$ by OLS. We repeat this many times and compute bias, mean-absolute error (MAE), and mean-squared error (MSE) for this idealized OLS estimator. We name it the “oracle” estimator since it is obtained in the absence of endogeneity and it is efficient by the Gauss-Markov Theorem.
@@
The performance of the oracle estimator is compared to that of the corresponding OLS event study estimator when we allow for correlation between $D_{t}$ and $\widetilde{u}_{t}$. In the latter regression, we successively increase the variance of $D_{t}$ and record how much the performance of the event study estimator approaches that of the oracle estimator in the idealized regression. In particular, we determine the threshold value for the variance of $D_{t}$ that ensures that the event study estimator performs as well as the oracle. Lastly, we verify whether the sample variance of $D_{t}$ in Nakamura and Steinsson (2018a) is larger than this threshold. If it is, relative exogeneity likely holds and we can interpret $\sigma_{e,t}^{2}$ as being large enough relative to the variance of $\widetilde{u}_{t}$ for the OLS event study estimator to have low enough bias to perform well in practice.
@@
Let $T_{C}$ denote the sample size of the control sample $t\in\mathbf{C}$. The policy sample consists of all regularly scheduled FOMC meeting days from 1/1/2000 to 3/19/2014. The control sample includes all Tuesdays and Wednesdays that are not FOMC meeting days over the same period. This yields $T_{P}=74$ and $T_{C}=762$ when the dependent variable is the 2-Year real yield, and $T_{P}=106$ and $T_{C}=1130$ when it is the 5-Year real yield.
@@
    Estimate the mean, variance and autoregressive coefficient of order 1 from all $D_{t}$ in the control sample ($t\in\mathbf{C}$). Denote these by $\overline{D}_{C}$, $\widehat{\sigma}_{D,C}^{2}$ and $\widehat{\rho}_{D,C,1}$. Similarly, obtain $\widehat{\sigma}_{Y,C}^{2}$ and $\widehat{\rho}_{Y,C,1}$ from all $Y_{t}$ in the control sample ($t\in\mathbf{C}$). Test whether $\widehat{\rho}_{D,C,1}$ and $\widehat{\rho}_{Y,C,1}$ are statistically significant different from zero.
@@
    Obtain the sample variances, $\widehat{\sigma}_{D,P}^{2}$ and $\widehat{\sigma}_{Y,P}^{2}$, from all $D_{t}$ and $Y_{t}$ in the policy sample ($t\in\mathbf{P}$).
@@
    If $Y_{t}$ is constructed as a change over a substantially longer frequency than $D_{t}$, compute $\widehat{\sigma}_{\widetilde{u}}^{2}=\widehat{\sigma}_{Y,P}^{2}\widehat{\sigma}_{D,C}^{2}/\widehat{\sigma}_{D,P}^{2}$. Note that $\widehat{\sigma}_{\widetilde{u}}^{2}$ is a proxy for the variance of $Y_{t}$ in the control sample. Since $Y_{t}$ is the 1-day change in the forward rate while $D_{t}$ is the change in a 30-minute window, the variance of $D_{t}$ is substantially smaller than that of $Y_{t}$ in the control sample and the variance of $Y_{t}$ and $\widetilde{u}_{t}$ are very similar. So we choose $\widehat{\sigma}_{\widetilde{u}}^{2}$ such that the ratio of the variances of $Y_{t}$ in the policy and control sample is the same as that of $D_{t}$, i.e., $\widehat{\sigma}_{Y,P}^{2}/\widehat{\sigma}_{\widetilde{u}}^{2}=\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}.$ [^19] Note that $\widehat{\sigma}_{\widetilde{u}}^{2}$ approximates the variance of $\widetilde{u}_{t}$ for $t\in\mathbf{P}$ when $D_{t}$ and $\widetilde{u}_{t}$ are independent and the variance of $\widetilde{u}_{t}$ is similar across the policy and control samples. Otherwise, compute $\widehat{\sigma}_{\widetilde{u}}^{2}=\widehat{\sigma}_{Y,C}^{2}-\beta^{2}\widehat{\sigma}_{D,C}^{2}$ with $\beta$ set equal to 0.99, which is the OLS estimate in Nakamura and Steinsson (2018a). [^20] We could choose any other value for $\beta$, the intuition behind the procedure and the results would not change.
  4. 4.
@@
    For $t=1,\ldots,\,T_{P}$ generate $\widetilde{u}_{t}\sim i.i.d.\,\mathscr{N}(0,\,\widehat{\sigma}_{\widetilde{u}}^{2})$ if $\widehat{\rho}_{Y,C,1}$ is not statistically different from zero. Otherwise generate $\widetilde{u}_{t}$ according to $\widetilde{u}_{t}=\widehat{\rho}_{Y,C,1}\widetilde{u}_{t-1}+\widetilde{v}_{t}$ where $\widetilde{v}_{t}\sim i.i.d.\,\mathscr{N}(0,\,(1-\widehat{\rho}_{Y,C,1}^{2})\widehat{\sigma}_{\widetilde{u}}^{2})$ and $\widetilde{u}_{0}=0$.
@@
    Generate $\widetilde{Y}_{t}=\beta\widetilde{D}_{t}+\widetilde{u}_{t}$ for $t=1,\ldots,\,T_{P}$ where $\widetilde{D}_{t}$ is generated in 4., $\widetilde{u}_{t}$ in 5. and $\beta$ is set equal to 0.99 as explained in 3. Run OLS for the regression of $\widetilde{Y}_{t}$ on $\widetilde{D}_{t}$ to obtain the OLS estimator $\widehat{\beta}_{\mathrm{orcale}}$.
  7. 7.
    Repeat 4-6. 5,000 times and compute the bias, MAE and MSE of $\widehat{\beta}_{\mathrm{orcale}}$.
  8. 8.
@@
    Find the value of $\delta$ such that $\mathrm{MSE}(\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta\right))=\mathrm{MSE}(\widehat{\beta}_{\mathrm{orcale}})$ [or $\mathrm{MAE}(\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta\right))=\mathrm{MAE}(\widehat{\beta}_{\mathrm{orcale}})$] and label it $\delta^{*}\left(\rho\right)$. Define $\widehat{\sigma}_{D,*}^{2}\left(\rho\right)=\widehat{\sigma}_{D,C}^{2}\left(1+\delta^{*}\left(\rho\right)\right).$
  11. 11.
    Repeat 8.-10. for a grid of $\rho$ values in [1, 1]11[-1,\,1][ - 1 , 1 ].
@@
Steps 1.-11. provide a detailed sensitivity analysis that computes the values of the variance of $D_{t}$ that allows the event study estimator to perform as well as the oracle estimator over a range of values of endogeneity $\rho$, namely $\widehat{\sigma}_{D,*}^{2}\left(\rho\right)$. If $\widehat{\sigma}_{D,P}^{2}\geq\widehat{\sigma}_{D,*}^{2}\left(\rho\right)$ one can be confident that the variance of the policy shock in Nakamura and Steinsson (2018a) is relatively large enough for a degree of endogeneity that satisfies $|\mathrm{Corr}\left(\widetilde{u}_{t},\,D_{t}\right)|\leq|\rho|$ for the event study estimator to have low enough bias for its MSE (MAE) to be as low as in the oracle regression with no endogeneity. If one is not interested in the full sensitivity analysis, but rather just in evaluating the values of $\rho$ for which the policy shock in Nakamura and Steinsson (2018a) is large enough to produce a credible event study estimator, one could replace $1+\delta$ in 9. by $\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}$ and instead of 10., find the value of $\rho$ such that $\mathrm{MSE}(\widehat{\beta}_{\mathrm{ES}}(\rho,\,\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}-1))=\mathrm{MSE}(\widehat{\beta}_{\mathrm{orcale}})$ [or $\mathrm{MAE}(\widehat{\beta}_{\mathrm{ES}}(\rho,\,\widehat{\sigma}_{D,P}^{2}/\widehat{\sigma}_{D,C}^{2}-1))=\mathrm{MAE}(\widehat{\beta}_{\mathrm{orcale}})$].
@@
The validity of the sensitivity analysis relies on the following two properties: (i) the model in (4.1) is correct also in the control sample, i.e., for $t\in\mathbf{C}$ and (ii) the order of magnitude of $\widehat{\sigma}_{\widetilde{u}}^{2}$ is a good proxy for the order of magnitude of $T_{P}^{-1}\sum_{t\in\mathbf{P}}\sigma_{\widetilde{u},t}^{2}$. Property (i) is implicitly assumed in typical VAR estimation of simultaneous equations models since typically the relationship between monetary policy and real economic variables is estimated with VARs at monthly, quarterly, yearly frequencies and each observation aggregates both FOMC announcement and non-announcement days. Property (ii) is reasonable since $\widetilde{u}_{t}$ for $t\in\mathbf{P}$ represents news and latent factors that are not specific to the announcement and would be present even if there were no FOMC announcements. Thus, its variation is expected to be similar to that in non-announcement days which in turn has the same order of magnitude as that of $Y_{t}$ in non-announcement days.
@@
To provide numerical support for the theoretical results of Theorem 1-4, we look at the change in the bias, MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}\left(\rho,\,\delta\right)$ as $\delta$ increases for a given $\rho\neq 0$ when looking at the 2-Year real yields. The results show that these summary statistics decrease quickly as $\delta$ grows. For example, for a small degree of endogeneity $\rho=0.25$, it is sufficient to increase the variance of the policy variable in the announcement windows nine-fold $\left(\delta=8\right)$ for the event study estimator to exhibit an MSE that is as small as that of the oracle estimator. A similar feature applies to MAE, though a slightly larger $\delta$ is needed. As we raise $\delta$ further, the MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}$ become smaller and smaller relative to those of $\widehat{\beta}_{\mathrm{oracle}}$. In particular, both the MAE and MSE of $\widehat{\beta}_{\mathrm{ES}}$ converge to zero quickly as $\delta$ increases for $\rho=0.25$.